The energy flux of sunlight reaching the surface of the earth is 1.388 ten poer three w/m2 .how many photons per square meter are incident on the earth per second assume that the photon in the sunlight have an average wavelength of 550nm.

Grade:12

1 Answers

Arun
5 years ago

Energy flux of sunlight reaching the surface of earth, Φ = 1.388 × 103 W/m2

Hence, power of sunlight per square metre, P = 1.388 × 103 W

Speed of light, c = 3 × 108 m/s

Planck’s constant, h = 6.626 × 10−34 Js

Average wavelength of photons present in sunlight, λ = 550 nm = 550 x 10-9 m

Number of photons per square metre incident on earth per second = n Hence, the equation for power can be written as: P= nE

∴ n = P/E = Pλ / hc = 1.388 x 103 x 550 x 10-9 / 6.626 x 10-34 x 3 x 108 = 3.84 x 1021 photons/m2/s

Therefore, every second, 3.84 x 1021 photons are incident per square metre on earth.